5) What is the standard deviation for the following set of values? 345mA 350mA 359mA348mA 351mA
Answer #1

Standard deviation is given by:

(a-T)

Where

x_bar = mean of given values

Now mean of given values will be

(All values are in mA)

x_bar = (345 + 350 + 359 + 348 + 351)/5 = 350.6

x (x - x_bar)^2/n
345 (345 - 350.6)^2/5 = 6.272
350 (350 - 350.6)^2/5 = 0.072
359 (359 - 350.6)^2/5 = 14.112
348 (348 - 350.6)^2/5 = 1.352
351 (351 - 350.6)^2/5 = 0.032

Now

(x - T) = (6.272 + 0.072 + 14.112 + 1.352 + 0.032) = 21.84

So,

Standard deviation = sqrt (21.84) = 4.67 mA

Please Upvote.

Comment below if you have any query.